OpenStudy (anonymous):

Best answer gets a medal! Determine a quadratic function with the given information. Answer in the form y=a(x-h)^2+k Write the equation of a quadratic function whose roots are -3/7 and 2/5 (answer without fractions)

OpenStudy (anonymous):

use y = a (x - h)^2 + k and put in the vertex values for h and k so if vertex is (-1, 4) y = a(x-(-1))² +4 y = a(x+1)² +4 now put in the point for x&y to solve for a --> point is (-3, - 4) so -4 = a(-3+1)² +4 -8 = a(-2)² -8 = a(4) -2 = a so the equation is y = -2(x+1)² +4

OpenStudy (anonymous):

Hope this helps!